Gdub08:

If 2x2(5-x)(3x+2) = 0, then what is the sum of all of the possible values of x?

Gdub08:

13/2 ? I think

SnowyBreaks:

no its not

SnowyBreaks:

okay give me one second

SnowyBreaks:

so I got the answer 2.

SnowyBreaks:

-2

SnowyBreaks:

Your doing eighth grade math correct?

SnowyBreaks:

Recalculated and the answer should be 2 and 1/3

SnowyBreaks:

2 1/3

SnowyBreaks:

always multiply the 2 by everything in the parenthesies

SnowyBreaks:

like 2(5-x)(3x+2) 2x5=10, 2 times -x equals -2x 2 times 3x=6x and 2x2=4 Next your gonna add 10 and 4 and that gets you 14. Then add 2x+-2x+6x

SnowyBreaks:

And that will get you 6x still.

SnowyBreaks:

Then divided both 6x by -6x and 14 by -6x

SnowyBreaks:

I hope this is helpful

Gdub08:

@snowybreaks wrote:
I hope this is helpful
thank youuuu.

SnowyBreaks:

@gdub08 wrote:
@snowybreaks wrote:
I hope this is helpful
thank youuuu.
no problem

V4MPZSKII:

Im too dumb for this

Gdub08:

@v4mpzskii wrote:
Im too dumb for this
lol

jhonyy9:

\[2x^2(5-x)(3x-2)=0\] 2x^2 = 0 => x=0 5-x=0 => x=5 3x-2=0 => 3x=2 => x=2/3 and now the sum of all these values of x = 0+5+2/3 \[x = 5 + \frac{ 2 }{ 3 }\] \[x = \frac{ 3*5 +2 }{ 3 }\] \[x = \frac{ 17 }{ 3 }\]